OpenStudy (anonymous):

Your company may wish to release a new e-reader device. Based on data collected from various sources, your company has come up with the following regression equation for the profit of the new e-reader: Profit = $0.15 × number of e-readers sold – $28 Or, assuming x = the number of e-readers sold, this would be the same regression equation: Profit = 0.15x – 28 In this case, the values are given in thousands (i.e., the cost of making an individual e-reader will be $150 [0.15 × 1,000], with $28,000 [28 * 1,000] in constant charges).

OpenStudy (anonymous):

Based on the results of the graph and the profit equation provided, discuss the relationship between profits and number of e-readers produced. (Hint: Consider the slope and y-intercept.) If the company does not sell a single e-reader, how much is lost ? Mathematically, what is this value called in the equation? If the company sells 5,000 e-readers, how much will the company make (or lose)? If profit must equal 100 thousand, how many e-readers will your company need to sell? (Round up to the nearest e-reader.) If your company is hoping to break even, how many e-readers will need to be sold to accomplish this? (Round up to the nearest e-reader.)Please submit your assignment.

OpenStudy (anonymous):

Profit = 0.15x – 28 1.There is a positive relationship between the profit and the number of e-readers produced, because of the slope 0.15 which is positive. That means when there is a increase in the number of e-readers produced, there is a increase in profits. Also, the y-intercept is negative. This means that there are pre-production costs. When production is 0 (x=0), the profit is negative. This implies the cost. 2. y-intercept is -28. When 0 e-readers is sold, the company loses 28 000. Mathematically this number is called the y-intercept. 3. When company sells 5000 e-reader, x= 5000. Profit = 0.15(5000) -28 = 722. Profit= $722 000 To earn 100 000. Profit= 100 100=0.15x - 28 x= 854 4. To break even, profit=0 0=0.15x -28 x=187

OpenStudy (anonymous):

Thanks .